Show that f(1, 1, 1), (1, – 1, 0), (1, 1, 0)1 is a
basis of R3 . Find a dual basis to this basis.
1
Expert's answer
2020-03-24T10:44:14-0400

Let's consider "e_1=(1,1,1), e_2=(1,-1,0),e_3=(1,1,0)" .

 Let's show that "e_1,e_2,e_3"  is a basis of "R^3" :

"\alpha_1 e_1+\alpha_2 e_2+\alpha_3 e_3=0" ,

where "\alpha_1 =\alpha_2=\alpha_3=0" .

 Really

"\alpha_1 +\alpha_2+\alpha_3=0\\
\alpha_1 -\alpha_2+\alpha_3=0\\
\alpha_1 =0"

Then "\alpha_1 =\alpha_2=\alpha_3=0" .

So "e_1,e_2,e_3"  is a basis of "R^3".


 Find a dual basis to this basis.

We can find the biorthogonal (dual) basis "\{e^1,e^2,e^3\}" by formulas below:

"e^1=(\frac{e_2\times e_3}{V})^T,e^2=(\frac{e_3\times e_1}{V})^T,e^3=(\frac{e_1\times e_2}{V})^T,"

where "^T" denotes the transpose and "V=(e_1,e_2,e_3)=e_1\cdot(e_2\times e_3)"

is the volume of the parallelepiped formed by the basis vectors "e_1,e_2,e_3" .

In our case

"V=\begin{vmatrix}
 1 & 1&1 \\
 1 & -1&0\\
1&1&0
\end{vmatrix}=1\cdot(-1)\cdot0+1\cdot0\cdot1+\\
+1\cdot1\cdot1-1\cdot(-1)\cdot1-1\cdot1\cdot0-1\cdot1\cdot0=2\\
e_2\times e_3=(\begin{vmatrix}
 -1& 0 \\
 1 & 0
\end{vmatrix},-\begin{vmatrix}
 1& 0 \\
 1 & 0
\end{vmatrix},\begin{vmatrix}
 1& -1 \\
 1 & 1
\end{vmatrix})=\\
=(0,0,2)\\
e_3\times e_1=(\begin{vmatrix}
 1& 0 \\
 1 & 1
\end{vmatrix},-\begin{vmatrix}
 1& 0 \\
 1 & 1
\end{vmatrix},\begin{vmatrix}
 1& 1 \\
 1 & 1
\end{vmatrix})=\\
=(1,-1,0)\\
e_1\times e_2=(\begin{vmatrix}
 1& 1 \\
 -1 & 0
\end{vmatrix},-\begin{vmatrix}
 1& 1 \\
 1 & 0
\end{vmatrix},\begin{vmatrix}
 1& 1 \\
 1 & -1
\end{vmatrix})=\\
=(1,1,-2)"

Then

"e^1=(\frac{(0,0,2)}{2})^T=\begin{pmatrix}
 0 \\
 0\\1
\end{pmatrix},\\
e^2=(\frac{(1,-1,0)}{2})^T=\begin{pmatrix}
 \frac{1}{2} \\
 -\frac{1}{2}\\0
\end{pmatrix},\\
e^3=(\frac{(1,1,-2)}{2})^T=\begin{pmatrix}
 \frac{1}{2}\\ \frac{1}{2} \\-1
\end{pmatrix}."

Denoting the indexed vector sets as "{\displaystyle B=\{v_{i}\}_{i\in I}}"  and "{\displaystyle B^{*}=\{v^{i}\}_{i\in I}}" , being biorthogonal means that the elements pair to have an inner product equal to 1 if the indexes are equal, and equal to 0 otherwise. Symbolically, evaluating a dual vector in V on a vector in the original space V:

"{\displaystyle v^{i}\cdot v_{j}=\delta _{j}^{i}={\begin{cases}1&{\text{if }}i=j\\0&{\text{if }}i\neq j{\text{,}}\end{cases}}}"

where "{\displaystyle \delta _{j}^{i}}"  is the Kronecker delta symbol.

In our case

"e^1\cdot e_1=0+0+1=1\\
e^1\cdot e_2=0+0+0=0\\
e^1\cdot e_3=0+0+0=0\\
e^2\cdot e_1=\frac{1}{2}-\frac{1}{2}+0=0\\
e^2\cdot e_2=\frac{1}{2}+\frac{1}{2}+0=1\\
e^2\cdot e_3=\frac{1}{2}-\frac{1}{2}+0=0\\
e^3\cdot e_1=\frac{1}{2}+\frac{1}{2}-1=0\\
e^3\cdot e_2=\frac{1}{2}-\frac{1}{2}+0=0\\
e^3\cdot e_3=\frac{1}{2}+\frac{1}{2}+0=1"

https://en.wikipedia.org/wiki/Dual_basis